What is the wavelength of light in nanometers that corresponds to the Balmer line from an n=12 transition? (Balmer lines end at an n=2.)
Answer #1

\lambda = B*(n2/(n2-m2))

\lambda is the wavelength.

B is a constant with the value of 3.6450682×10−7 m or 364.50682 nm.

m is equal to 2

n is an integer such that n > m.here n = 12

there fore \lambda = 3.6450682×10−7×(122/(122-22))

\lambda = 3.6450682×10-7×(144/(144-4))

\lambda = 3.6450682×10-7×1.02

\lambda = 3.7492×10-7

\lambda = 0.37492×10-8

\lambda = 0.37492 Ao